# Understanding oscillating features of the time-like nucleon electromagnetic form factors within the extending vector meson dominance model

### 08 Dec 2023 v1

@@ III NUMERICAL RESULTS
Figure 4: The ratio $R=\sigma_{e^{+}e^{-}\to p\bar{p}}/\sigma_{e^{+}e^{-}\to n\bar{n}}$ as a function of $\sqrt{s}$ compared with the experimental data.
@@ IV SUMMARY
Our study here shows that the so-called periodic behaviour of the nucleon effective form factors is not confirmed. However, there are indeed nonmonotonic structures in the line shape of nucleon effective form factors, which can be naturally reproduced by considering the contributions from the low-lying excited vector states. We hope that our present work can stimulate more studies along this line, and it is expected that more precise future experimental measurements, especially for the case of the $e^{+}e^{-}\to n\bar{n}$ reaction, can be used to check our model calculations.
@@ ACKNOWLEDGEMENTS
We would like to thank Dr. Lei Xia for useful discussions. This work is partly supported by the National Natural Science Foundation of China under Grant No. 12075288. It is also supported by the Youth Innovation Promotion Association CAS.

### 29 Feb 2024 v2

@@ III NUMERICAL RESULTS
Figure 6: The ratio $R=\sigma_{e^{+}e^{-}\to p\bar{p}}/\sigma_{e^{+}e^{-}\to n\bar{n}}$ as a function of $\sqrt{s}$ compared with the experimental data which are taken from: FENICE 1998 Antonelli et al. (1998) and BESIII 2021 Ablikim et al. (2021b).
Moreover, we also calculated the relative phase $\Delta\Phi$ between $G_{E}$ and $G_{M}$ of the nucleon. The relative phase $\Delta\Phi$ is related to the spin polarization of baryons in the $e^{+}e^{-}\to B\bar{B}$ reaction Fäldt and Kupsc (2017); Ablikim et al. (2020d, 2022). Here, we can write
$G_{E}/G_{M}=e^{i\Delta\Phi}|G_{E}/G_{M}|.$ (16)
Using the fitted parameters in Table 2, we can obtain the $\Delta\Phi$ as a function of $\sqrt{s}$ for both proton and neutron, which are shown in Fig. 7. In the reaction threshold, the relative phase $\Delta\Phi$ equals 0, since the electric and magnetic form factors are equivalent. As the center of mass energy $\sqrt{s}$ increases, the value of $\Delta\Phi$ undergoes significant changes, particularly in the vicinity of $\sqrt{s}=2.30$ GeV. This is because the value of $G_{E}$ for proton is close to zero and it is very small for the neutron. Since the strength of the polarization is proportional to the $\sin\Delta\Phi$, according to the numerical results as shown in Fig. 7, it implies that there is obvious spin polarization at 2.28 GeV and 2.40 GeV, and the spin polarization has opposite directions at the two energy points. It is expected that the future experiments can be used to check our model calculations.
Figure 7: The predictions of $\Delta\Phi$ for proton and neutron as a function of $\sqrt{s}$. The error bands are obtained with the uncertainties of fitted parameters.
@@ IV SUMMARY
Our study here shows that the so-called periodic behaviour of the nucleon effective form factors is not confirmed. However, there are indeed nonmonotonic structures in the line shape of nucleon effective form factors, which can be naturally reproduced by considering the contributions from the low-lying excited vector states. We hope that our present work can stimulate more studies along this line, and it is expected that more precise experimental measurements in near future, especially for the case of the $e^{+}e^{-}\to n\bar{n}$ reaction, can be used to check our model calculations.
@@ ACKNOWLEDGEMENTS
We would like to thank Ji-Feng Hu and Lei Xia for useful discussions. This work is partly supported by the National Natural Science Foundation of China under Grant No. 12075288. It is also supported by the Youth Innovation Promotion Association CAS.